what is the answer of this equation 3/8-1/4=1/2t-3/4
Answer by oberobic(2304)

3/8 -1/4 = 1/2t -3/4
.
add 3/4 to both sides
.
3/8 -1/4 + 3/4 = 1/2t
.
1/2t = 3/8 + 2/4
.
1/2t = 3/8 + 4/8
.
1/2t = 7/8
.
cross multiply
.
8t = 14
.
t = 14/8
.
t = 7/4
.
Done.